### v2

# Think-at-Hard: Selective Latent Iterations to Improve Reasoning Language Models

Abstract
Improving reasoning abilities of Large Language Models (LLMs), especially under parameter constraints, is crucial for real-world applications. Looped transformers address this by performing multiple latent iterations to refine each token beyond a single forward pass. However, we identify a latent overthinking phenomenon: most token predictions are already correct after the first pass, but are sometimes revised into errors in later iterations. In this work, we ask whether selectively skipping latent iterations may improve accuracy. We reveal significant potential with an oracle iteration policy that boosts model performance by up to 7.3%. Motivated by this, we propose Think-at-Hard (TaH), a looped transformer optimized for selective iteration. TaH employs a lightweight neural decider to trigger latent iteration only at tokens that are likely incorrect after the standard forward pass. During latent iterations, depth-aware Low-Rank Adaptation (LoRA) modules shift the LLM’s objective from general next-token prediction to focused hard-token refinement. A duo-causal attention mechanism extends attention from the token sequence dimension to an additional iteration depth dimension, enabling cross-iteration information flow with full sequential parallelism. Experiments on nine benchmarks show consistent gains across math, QA, and coding tasks. With identical parameter counts, TaH outperforms always-iterate baselines by 3.8–4.4% while skipping iterations on 93% of tokens, and exceeds single-iteration Qwen3 baselines by 3.0–3.8%. When allowing $3\%$ more parameters from LoRA and decider modules, the gains further increase to 5.3–6.2% and 6.1–6.8%, respectively. Our code is available at https://github.com/thu-nics/TaH.
1 Introduction
Figure 1: Latent iterations can fix wrong predictions, but can also overthink and flip correct ones. Selective iteration only when needed can improve reasoning with reduced computation.
Recent advances in Large Language Model (LLM) reasoning have enabled broad applications across diverse domains (Jaech et al., 2024; Guo et al., 2025; Yang et al., 2025). With hundreds of billions of parameters, LLMs can generate complex Chain-of-Thought (CoT) to solve challenging tasks. At the same time, smaller language models have also drawn increasing attention. With only a few billion parameters, they offer compelling alternatives: lower costs, faster inference, and suitability for edge computing (Abdin et al., 2024; Team et al., 2025; Wang et al., 2025a).
At this crossroads, enhancing reasoning capabilities under parameter constraints becomes a central challenge. A common approach is to distill smaller models to mimic LLM CoT trajectories. However, not all tokens are equally predictable: certain tokens encode critical logic or reasoning directions that are fundamentally harder to predict (Lin et al., 2024; Fu et al., 2025; Wang et al., 2025b). With limited computation per output token, small models quickly hit a performance ceiling and mispredict some of these tokens. Once critical errors occur, the reasoning trajectory can irrecoverably diverge and produce drastically different outcomes.
Prior work proposes looped transformers to address this parameter–performance paradox (Hutchins et al., 2022; Saunshi et al., 2025; Zeng et al., 2025; Zhu et al., 2025). Instead of verbalizing the next token immediately after one forward pass, these models typically feed the last-layer hidden states back into the LLM for additional passes, refining representations in the latent space. After certain iterations, the final hidden states pass through the language modeling head to generate the next token. By uniformly scaling up iterations per token, these models can correct initially wrong token predictions, potentially increasing performance without increasing parameter count.
However, we identify a latent overthinking problem in looped transformers, where excessive iterations revise correct answers into wrong ones. As shown in Figure 3, while the second iteration corrects 8.7% of predictions, it also flips 2.1% of correct ones into errors. This occurs because most tokens, such as coherence or suffix tokens, are already predicted correctly after the first pass; further iterations may instead introduce harmful changes. This mirrors overthinking in explicit CoT reasoning (Wu et al., 2025), where additional reasoning steps degrade rather than improve answers. This reveals a surprising opportunity:
selectively skipping latent iterations on most tokens can further increase model performance.
We validate this with an oracle policy that iterates only on initially mispredicted tokens, as shown in Table 1. Compared to always iterating, this selective oracle can achieve up to 32% higher accuracy with optimized architecture.
Achieving selective latent iteration presents three main challenges. First, the model architecture should enable cross-depth attention, allowing each iteration to access full context. This is crucial because when early tokens skip deeper iterations, later tokens must still access their representations from shallower depths. Meanwhile, this cross-depth flow cannot compromise the sequence-level parallelism essential for efficient training and prefilling. Second, the model must adapt to distribution shifts across iterations, while maximizing parameter reuse. Third, training must remain stable despite tightly coupled dependencies: the iteration policy depends on prediction quality at each depth, while that quality depends on the depths to which the policy assigns previous tokens.
Table 1: Performance comparison across iteration strategies, using Ouro-1.7B model except for the last row. Always1 means no latent iteration. Oracle policy iterates on 12–19% tokens. MMLU100 denotes the first 100 questions from MMLU-STEM.
Iter. Policy NTP AMC23 MMLU100 HE++
Always1 73.1 38.1 56.0 39.6
Always2 79.7 40.6 60.0 40.9
Orcl. 81.8/+2.1/+2.1 47.9/+7.3/+7.3 62.0/+2.0/+2.0 43.3/+2.4/+2.4
Orcl. w.TaH 89.3/+9.6/+9.6 68.8/+28.2/+28.2 85.0/+25.0/+25.0 72.9/+32.0/+32.0
To address these challenges, we propose TaH, a looped transformer optimized for selective latent iteration. As shown in Figure 2, TaH employs a neural decider to determine whether to iterate or verbalize each token. We design a duo-causal attention mechanism to enable cross-depth attention with full sequence parallelism. To specialize deeper iterations for current-token refinement and preserve strong first-pass predictions, we apply LoRA adapters solely at iterations $d>1$. We train TaH stably by aligning both the LLM backbone and iteration decider with a static oracle iteration policy. We summarize our contributions as follows.
  • Selective Latent Iteration. We identify the latent overthinking phenomenon and quantify its influence on token prediction accuracy and downstream tasks. This insight motivates new directions where latent iteration is applied selectively to a few tokens for both better reasoning quality and efficiency.
  • Specialized Model Architecture. We develop a model architecture that natively supports selective iterations. The dedicated duo-causal attention mechanism, LoRA adapters, and iteration decider enable efficient cross-depth information flow, objective transitions, and dynamic depth selection.
  • Stable Training. We introduce a stable training scheme that uses a static oracle policy to decouple model adaptation and policy learning. It overcomes the circular dependency between iteration decisions and prediction quality.
We fine-tune TaH from Qwen3-Base 0.6B and 1.7B on Open-R1, then tested across nine reasoning benchmarks spanning math, QA, and coding tasks. TaH achieves average accuracy gains of 3.0% and 3.8% over standard single-iteration baselines. Compared to Ouro looped-transformer trained on the same data, TaH achieves 3.8–4.4% gains while reducing latent iterations by 93%.
2 Related Work
Unlike standard LLMs that verbalize at every autoregressive step, latent thinking shifts part of generation away from explicit natural-language CoT in order to improve reasoning (Li et al., 2025). We provide additional discussions on signal-guided control and latent optimization in Appendix A.5.
Looped Transformers. These methods interleave latent and verbal reasoning, adding latent iterations before each token verbalization. Previous work focuses on scaling up iteration depths, with the main architectural focus on next-iteration inputs: reusing hidden states directly (Saunshi et al., 2025; Geiping et al., 2025; Zhu et al., 2025) or using logit-weighted embeddings (Zeng et al., 2025). Looped transformers achieve deeper computation without parameter increases. However, uniform depth scaling burdens training and inference, and risks overthinking already-correct tokens.
Positioning. TaH belongs to the looped transformer family, but identifies selective iteration as a new design principle to improve performance. While concurrent works (Bae et al., 2025; Zhu et al., 2025) also enable dynamic recursion, they degrade performance at non-maximum depths and require full retraining. TaH instead leverages existing pre-trained models, adding depth-aware LoRA and duo-causal attention to improve reasoning with minimal fine-tuning overhead.
3 Preliminaries
Autoregressive LLMs. Modern LLMs generate text through an autoregressive next-token prediction process. It includes a prefill stage and a decode stage (Radford et al., 2018; 2019; Kwon et al., 2023). In the prefill stage, the model processes the entire input sequence in parallel; in the decode stage, it consumes one new token at a time along with cached history to predict the next token.
Formally, let $t_{i}$ denote the token at position $i$ and $x_{i}\in\mathbb{R}^{h}$ its embedding. Let $E\in\mathbb{R}^{v\times h}$ be the embedding matrix, so $x_{i}=E[t_{i}]$ when $t_{i}$ is treated as an index. Here, $v$ and $h$ are the vocabulary size and hidden dimension. The output projection matrix is $W_{\text{out}}\in\mathbb{R}^{h\times v}$ (equal to $E^{\top}$ if tied). Given the context $T_{\leq i}=[t_{0},\dots,t_{i}]$ with embeddings $X_{\leq i}=[x_{0},\dots,x_{i}]$, the model $\theta$ produces a last-layer hidden state $y_{i}$ for token $t_{i}$:
$y_{i}\;=\;\mathcal{P}_{\theta}\!\big(x_{i}\mid X_{\leq i}\big)\;\in\;\mathbb{R}^{h}.$ (1)
The next-token distribution $p_{i}$ and sample are:
$p_{i}\;=\;\mathrm{softmax}\!\big(W_{\text{out}}^{\top}y_{i}\big)\in\mathbb{R}^{v},\qquad t_{i+1}\;=\;\mathcal{S}(p_{i}),$ (2)
where $\mathcal{S}$ is a sampling rule such as nucleus sampling. Decoding repeats until an end-of-sequence token is generated.
Causal Attention. Modern LLMs typically adopt causal attention mechanism. As shown in Figure 2(a), each position attends only to itself and earlier positions, consistent with Equation 1. This design brings two key benefits: (1) it enables parallel training with next-token prediction and shifted logits, avoiding the need for token-by-token generation; and (2) it allows efficient inference by caching Key/Value states of past tokens instead of recomputing them.
Looped Transformers. Looped transformers introduce an inner loop that iterates in latent space before verbalizing each output token. Let $d\in\{1,2,\dots\}$ denote the iteration depth (written as a superscript), and set $x_{i}^{(0)}=E[t_{i}]$. At each iteration, looped transformers update $y_{i}$ with causal attention on the hidden states of the current iteration:
$y_{i}^{(d)}\;=\;\mathcal{P}_{\theta}\!\big(x_{i}^{(d)}\,\big|\,X_{\leq i}^{(d)}\big),\qquad X_{\leq i}^{(d)}=[x_{0}^{(d)},\dots,x_{i}^{(d)}].$ (3)
An inner transition then produces the next-depth embedding. For example, Loop (Saunshi et al., 2025) simply sets $x_{i}^{(d+1)}=y_{i}^{(d)}$, while Ponder (Zeng et al., 2025) uses a logit-weighted embeddings:
$x_{i}^{(d+1)}=\mathrm{softmax}\!\big(W_{\text{out}}^{\top}y_{i}^{(d)}\big)\ E=p_{i}^{(d)}E.$ (4)
In practice, it uses the top-100 logits instead of full logits for efficiency.
Verbalization occurs at a fixed maximum depth $d_{\max}$ shared by all tokens, where $y_{i}^{(d_{\max})}$ is transformed into the next token $t_{i+1}$, resembling Equation 2.
4 Selective Latent Iteration Oracles
Setup. We analyze the latent iteration behavior of the Ouro looped transformer (Zhu et al., 2025) and our proposed TaH model (architecture detailed in Section 5). All models are fine-tuned from Qwen3-1.7B-Base on a balanced 100K-sample subset of the Open-R1 dataset (Hugging Face, 2025), following setups in Section 6.1.
Oracle Iteration Policy. To investigate the potential of selective iteration, we establish an oracle policy $\pi$. It triggers additional iterations only when the reference LLM $\theta$ mispredicts the target token at the first forward pass. In this section, we set $\theta$ to the model under evaluation, so $\pi$ acts as a greedy, locally optimal policy.
Formally, let $\hat{p}_{i}$ denote the reference model’s first-pass next-token distribution at position $i$, and $t_{i+1}$ the ground-truth token. The oracle iteration depth $d_{i}^{\pi}$ is:
$d_{i}^{\pi}\;=\;1+\mathcal{D}(\hat{p}_{i}^{(1)},t_{i+1}),$ (5)
where $\mathcal{D}$ is a binary discrepancy metric. We use top-1 mismatch here: $\mathcal{D}=\bm{1}[\hat{t}_{i+1}\neq t_{i+1}]$, with $\hat{t}_{i+1}=\arg\max_{t}\hat{p}_{i}^{(1)}$ ($\bm{1}[\cdot]$ denotes the indicator function). For simplicity, we assume $d_{\max}=2$. The arbitrary-depth case is in Appendix A.2.5, and alternative discrepancy metrics are ablated in Table 13.
For brevity, we call a token easy if the reference model correctly predicts it at the first pass ($d_{i}^{\pi}=1$), and hard otherwise. Prior work (Fu et al., 2025) shows that hard tokens typically occupy only a small proportion (e.g., $7\%$) of all tokens.
Next-Token Prediction. We evaluate next-token prediction (NTP) accuracy by comparing top-1 predictions against ground-truth tokens in the Open-R1 validation set. As shown in Table 1, always iterating twice improves NTP accuracy for Ouro by 6.6%, but at the cost of doubled depth. Surprisingly, the oracle policy, which skips iterations on 81-88% tokens, further improves accuracy by 2.1%. This gain comes from avoiding latent overthinking: without selective iteration, the model can revise correct predictions to wrong ones at the second pass, as shown in Figure 3.
Downstream Tasks. We next examine whether NTP gains translate to downstream task performance. Since ground-truth tokens are unavailable during generation, we use top-1 predictions from the stronger Qwen3-8B as proxy labels. Table 1 shows that the oracle policy improves downstream performance by 2.0–7.3%.
TaH Design Objectives. The oracle experiments reveal two key insights. First, models have significant untapped potential when iteration depth is selected correctly. Second, the oracle policy is effective enough to learn from, even though other globally optimal policies may exist. These findings motivate the design objectives of TaH: (1) its model architecture should natively support selective iteration depths; and (2) following the oracle policy, its training objective for deeper iterations is not to cover all tokens, but to selectively focus only on the few failing tokens. We later validate that TaH-1.7B can better utilize the oracle policy to achieve $25\%$ improvement, surpassing even Qwen3-4B. While the oracle requires ground-truth tokens unavailable at inference, approximating it with neural networks is a promising direction.
5 TaH Design
Figure 2: TaH Overview. (a) Regular causal attention: tokens attend only to previous positions. (b) Our duo-causal attention: tokens attend to both previous positions and shallower iteration depths, maintaining 2D causality. (c) Model architecture: TaH selectively iterates or verbalizes tokens. It uses LoRA at deeper iterations to shift from next-token prediction to hard-token refinement. A neural decider determines whether to continue iterating or output the token.
We expand the motivations and key designs of TaH in this section, including the duo-causal attention mechanism (Section 5.1), model architecture (Section 5.2), and training scheme (Section 5.3).
5.1 Duo-Causal Attention
Motivation. In looped transformers with fixed depth, standard causal attention on the current iteration’s KV states incorporates all context (Equation 3). However, dynamic iteration depths create a challenge: tokens at deeper levels cannot access hidden states of previous tokens that verbalized at shallower depths. This poses a dilemma between requiring up-to-date context from all previous tokens and maintaining parallel training where depth-$d$ computations cannot depend on uncomputed deeper states ($d^{\prime}>d$). Existing approaches compromise on one of these aspects. Some sacrifice parallelism by allowing attention to deeper iterations (Hao et al., 2024); others preserve parallelism by restricting attention to only the initial iteration’s KVs (Bae et al., 2025). To resolve this dilemma, we introduce a simple yet effective mechanism to maximize cross-depth information flow while maintaining high parallelism.
Duo-causal Attention Mechanism. As shown in Figure 2(b), duo-causal attention extends causality to two dimensions, letting tokens attend across both previous positions and shallower iteration depths. Formally, we extend the accessible set from Equation 3 to
$X_{\leq i}^{(\leq d)}\;=\;\{\,x_{j}^{(k)}\;|\;j\leq i,\;k\leq d\,\}.$ (6)
When all tokens iterate only once (as in standard transformers), this reduces to regular causal attention. The duo-causal design achieves both full parallel training and cross-depth information flow. At depth $d$, all tokens compute their depth-$d$ representations simultaneously using only and all information from depths $1$ through $d$.
5.2 Depth-adaptive Model Architecture
Motivation. Previous looped transformers typically use identical weights across all iterations. However, we find that over 73% of next-tokens are correctly predicted at the first iteration (Table 1). This suggests deeper iterations serve a different objective: they refine the first iteration’s prediction rather than predicting further ahead to the next-next token. This mirrors deep LLMs, where shallow layers predict next tokens for deeper layers to refine (Belrose et al., 2023; Schuster et al., 2022; Bae et al., 2023). While deep LLMs naturally handle this shift through distinct parameters per depth, looped transformers must accommodate both objectives with shared weights, potentially limiting performance. Moreover, fixed iteration depths can cause latent overthinking, motivating our dynamic approach.
Backbone Model. To address the objective shift, we apply a LoRA adapter (Hu et al., 2022) to the shared LLM backbone only for iterations $d>1$. As shown in Figure 2(c), this allows the base LLM to focus on latent embeddings, while the adapter handles the objective shift. It preserves strong next-token prediction at $d=1$, alleviating interference from deeper iterations. We also add residual connections across iterations to simplify the refinement and improve gradient flow. Formally, at depth $d$, we compute
$y_{i}^{(d)}\;=\;\mathcal{P}_{\theta_{d}}\!\Big(x_{i}^{(d)}\,\Big|\,X_{\leq i}^{(\leq d)}\Big),$ (7)
with depth-specific parameters
$\theta_{d}=\theta\ \text{for }d=1,\qquad\theta_{d}=\theta+\Delta\ \text{for }d>1,$
where $\theta$ and $\Delta$ denote the LLM and LoRA weights, respectively. The next-iteration inputs use logit-weighted embeddings (Equation 4); verbalization follows standard sampling (Equation 2). Each $y_{i}^{(d)}$ either continues iterating or verbalizes according to the decider $\mathcal{I}_{\phi}$.
Iteration Decider. We use a lightweight MLP as the iteration decider $\mathcal{I}_{\phi}$ to determine whether each token should continue iterating or verbalize. After each iteration, it processes concatenated hidden states from shallow, middle, and final layers of the backbone LLM to predict a continuation probability:
$\hat{c}_{i}^{(d)}\;=\;\mathcal{I}_{\phi}\!\big(h_{i}^{(d)}\big)\in[0,1].$
During inference, token $i$ verbalizes when $\hat{c}_{i}^{(d)}$ falls below threshold $c_{\text{threshold}}$ or reaches maximum depth $d_{\max}$.
5.3 Training Scheme
We adopt a two-stage training scheme. We first fine-tune the backbone model to support selective latent iteration, then train the iteration decider. Both stages are aligned to the same oracle iteration policy.
Motivation. As shown in Figure 2(c), the backbone LLM and the iteration decider are tightly coupled, making joint training unstable. Specifically, the backbone’s prediction quality across iterations determines the optimal depth, while the decider controls the backbone’s KV cache and output depth. To stabilize training, we train the two components sequentially under a fixed oracle policy $\pi$ (validated in Section 4).
Oracle Iteration Policy $\pi$.
Our goal is to iterate only on hard tokens that a standard supervised fine-tuned (SFT) model would mispredict on the first pass. Thus, we define $\pi$ with SFT model as the reference: we trigger an additional iteration when its top-1 next-token prediction differs from the ground-truth token. In principle, one could instead define $\pi$ using the current looped model itself (i.e., an on-policy oracle), but we find this to be empirically unstable (Section 6.3).
Stage 1: Backbone supervision under $\pi$.
We optimize the backbone LLM ($\theta$ and LoRA adapter $\Delta$) with $\pi$-guided iteration execution. The loss is standard next-token prediction at the oracle-determined depth:
$\mathcal{L}_{\text{SFT}}(\theta,\Delta)\;=\;\sum_{i}\;-\log p_{i}^{(d_{i}^{\pi})}\!\big(t_{i+1}\big),$
where $p_{i}^{(d_{i}^{\pi})}$ is the next-token distribution at position $i$, depth $d_{i}^{\pi}$. This preserves first-iteration accuracy for easy tokens while training deeper iterations to refine hard tokens.
Stage 2: Decider imitation under frozen backbone.
We freeze the backbone model $(\theta,\Delta)$ and train the iteration decider $\phi$ to imitate the oracle policy’s continuation decisions. We minimize weighted binary cross-entropy:
$\mathcal{L}_{\text{dec}}(\phi)\;=\;-\sum_{i,d}w_{d}\Big[c_{i}^{(d)}\log\hat{c}_{i}^{(d)}+(1-c_{i}^{(d)})\log(1-\hat{c}_{i}^{(d)})\Big],$
where the sum ranges over tokens $i$ and depths $d=1,\ldots,\min\{d_{\max}-1,d_{i}^{\pi}\}$. Here $c_{i}^{(d)}$ is the ground-truth continuation label, $\hat{c}_{i}^{(d)}$ is the predicted probability, and $w_{d}$ is the class weight for label imbalance (ratio of stop to continue labels).
The two-stage scheme stabilizes training by decoupling backbone learning (conditioned on a fixed $\pi$) from iteration policy learning (imitation of $\pi$).
6 Experiment
6.1 Setup
We present key experiment configurations here, with detailed setups in Appendix A.1.
Baselines. We compare diverse methods under equal parameter budgets, using Qwen3-{0.6B,1.7B,4B}-Base (Yang et al., 2025) as backbones. We compare TaH over the following baselines: (1) Standard, which always verbalizes directly and reduces to the standard Qwen model; (2) SoftThink, a latent optimization method, implemented following official design (Zhang et al., 2025) on top of the Standard model; (3) AlwaysThink, similar to TaH, but always iterates twice for all tokens during training and inference; (4) Ouro, a looped transformer that can scale iteration depths, implemented following official design (Zhu et al., 2025). Unless otherwise specified, TaH, Ouro, and AlwaysThink all use a maximum of two iterations, fine-tuned from the same Qwen3 backbone on the same training data and recipe. Ouro reaches its highest performance when set to maximum iterations, so we report results under this setup.
TaH Setup. To match the total parameter count of TaH (with LoRA and decider) with that of baselines, we prune one layer from the LLM backbone before training. The layer is chosen to minimize the increase in validation loss. We also report results for an unpruned variant, TaH+, which adds less than 3% extra parameters from LoRA and decider. The detailed parameter composition is shown in Table 4. Following (Fu et al., 2025), we set the continuation threshold $c_{\text{threshold}}=0.9$ with about 7% of tokens being iterated twice. The oracle policy $\pi$ uses Qwen3-0.6B, 1.7B, and 4B as reference models, respectively.
Training Scheme. All models are trained on the balanced Open-R1 (Hugging Face, 2025) mixture (math, QA, and code; 100K samples) using supervised fine-tuning. To fit memory and compute limits, we exclude responses longer than 8,192 tokens; 4B models additionally truncate at 4,096 tokens; all other training settings follow the official Open-R1 script. Each method is sufficiently trained for 5 epochs, and we select the checkpoint with the lowest validation loss as the final model. All backbones are initialized from the corresponding Qwen3-Base model.
Evaluation Setup. We evaluate across challenging reasoning benchmarks, including GSM8K (Cobbe et al., 2021), MATH500 (Hendrycks et al., 2021b), AMC23 (American Mathematics Competitions), AIME25 (American Invitational Mathematics Examination), OlympiadBench (He et al., 2024), MBPP++ (Austin et al., 2021), HumanEval++ (Chen et al., 2021), GPQA-Diamond (denoted as GPQA) (Rein et al., 2023), and MMLU-STEM  (Hendrycks et al., 2021a). The maximum generation length is set to 8,192 tokens for all benchmarks, except GSM8K and MMLU-STEM which use 4,096 due to their simpler problems and larger size. Performance is reported as pass@1 under a zero-shot CoT setting, using sampling temperature 0.6. We generate one sample per problem for large datasets (MATH500, OlympiadBench etc.), and eight samples per problem for small datasets (AMC23, AIME25).
6.2 Performance
Benchmark Evaluation. We validate TaH’s reasoning ability across all nine benchmarks. Table 5 presents performance results for models at 0.6B, 1.7B, and 4B parameter sizes. Compared with the strong Standard Qwen3 baselines, we observe that existing approaches (AlwaysThink, SoftThink, and Ouro) yield only marginal improvements when fine-tuned from base. In contrast, TaH achieves consistent gains across benchmarks. For 0.6B and 1.7B models, TaH delivers average improvements of 3.0% and 3.8% over Standard, respectively; TaH+, which adds less than 3% additional parameters, further pushes these gains to 5.3% and 6.2%. Compared to concurrent work Ouro, TaH and TaH+ achieve 3.8–4.4% and 6.1–6.8% gains, respectively. For 4B models, which are trained with 4K context length due to resource limits, TaH and TaH+ also achieve gains of 1.7% and 2.2%, with potential for further improvement at longer training contexts.
Table 2: Accuracy comparison across benchmarks. Best results are highlighted in bold. 4B models trained with $\leq$4K lengths due to resource constraints; AlwaysThink is excluded due to Out-Of-Memory (OOM) during training.
Standard SoftThink Ouro AlwaysThink TaH TaH+
0.6B
AIME25 1.9 2.9 2.1 1.3 2.1 4.6
OlympiadBench 15.4 14.0 14.2 12.6 19.1 20.6
AMC23 22.7 22.2 19.7 21.9 24.1 24.7
MATH500 39.9 39.6 37.4 37.8 46.2 51.8
GSM8K 58.2 55.9 56.6 52.6 63.6 67.6
GPQA 31.1 24.7 35.4 30.8 29.0 31.3
MMLU-STEM 54.2 53.0 54.0 51.4 56.4 59.0
HumanEval++ 16.8 14.3 18.9 9.1 21.6 22.0
MBPP++ 28.8 29.5 23.5 13.8 33.9 35.1
Average 29.9 28.5 29.1 25.7 32.9/+3.0/+3.0 35.2/+5.3/+5.3
1.7B
AIME25 10.8 5.4 10.8 7.5 13.8 15.4
OlympiadBench 33.8 30.7 31.3 31.0 37.2 37.6
AMC23 39.7 40.3 40.6 40.9 40.9 48.4
MATH500 67.8 64.8 68.2 63.2 71.4 72.6
GSM8K 80.2 80.0 79.8 74.2 84.8 84.5
GPQA 30.3 33.3 32.8 30.5 33.3 39.4
MMLU-STEM 74.1 73.5 72.4 69.6 74.8 76.6
HumanEval++ 39.0 43.3 40.9 16.4 50.0 51.5
MBPP++ 51.9 49.1 45.5 25.6 55.3 57.5
Average 47.5 46.7 46.9 39.9 51.3/+3.8/+3.8 53.7/+6.2/+6.2
4B
AIME25 24.2 25.8 25.0 27.1 27.9
OlympiadBench 50.4 50.2 51.4 50.5 52.6
AMC23 64.2 63.1 64.4 69.7 68.1
MATH500 84.2 85.0 83.8 85.8 85.6
GSM8K 91.7 92.5 90.7 91.0 91.7
GPQA 46.2 50.3 50.0 48.5 49.0
MMLU-STEM 85.4 86.0 85.9 86.2 86.6
HumanEval++ 69.2 69.2 70.7 70.1 72.0
MBPP++ 65.9 66.7 66.7 67.7 68.1
Average 64.6 65.4 65.4 66.3/+1.7/+1.7 66.8/+2.2/+2.2
Training Dynamics. During stage 1 (LLM backbone training), TaH performs iterations according to the oracle policy. As shown in Figure 4(a), it converges notably faster than the Standard baseline and also achieves much lower validation perplexity. During stage 2 (iteration-decider training), the neural decider successfully imitates the oracle strategy. It reaches about 83% accuracy at predicting the oracle’s iteration decisions, as shown in Figure 9.
6.3 Design Choice Exploration
We demonstrate the effectiveness and robustness of each design choice of TaH through ablation studies. All experiments in this section train TaH and its variants on the math subset of Open-R1 and evaluate on MATH500, AMC23, and OlympiadBench.
Model Architecture. (1) Iteration Scheme. As shown in Table 13, decider-based iteration outperforms the Always-1 and Always-2 alternatives, confirming the practical benefits of selective iteration, even with imperfect decisions from neural decider. Note that for Always-1, duo-causal attention degenerates to regular causal attention. (2) Duo-Causal Attention. Replacing duo-causal attention with standard causal attention variants causes significant drops: (a) attending only to the first iteration (Causal@iter1) drops by 5.4%; (b) attending only to the current iteration (Causal@current) drops even more, by 8.5%. They confirm duo-causal attention’s essential role in cross-depth information flow. (3) Depth Adapters. Removing LoRA and residual connections leads to consistent drops, confirming their beneficial roles in objective transition across iterations.
Training Scheme. (1) Supervision type. Some previous work supervises all iteration depths with next-token labels, while TaH only supervises final output tokens. As shown in Table 13, such token+latent supervision underperforms TaH. This aligns with our intuition that different iterations should focus on their respective objectives. (2) Iteration policy during LLM training. We compare our static oracle strategy $\pi$ with two alternatives. The decider-based approach trains the iteration decider first, then uses it during backbone training. It suffers from the coupling challenge discussed in Section 5.3. The dynamic approach recalculates the oracle using the evolving backbone in Equation 5, facing the same coupling challenge and causing training collapse. These results support our selection of the static oracle policy. (3) Discrepancy metric in $\pi$. We compare three discrepancy metrics to trigger latent iteration in $\pi$: top-1 mismatch, entropy, and cross-entropy (detailed definitions in Appendix A.3.1). As Table 13 shows, top-1 mismatch yields the best result, confirming its empirical effectiveness.
Continuation Threshold. As shown in Figure 4(b), TaH maintains robust performance across different continuation thresholds and iteration ratios. We empirically set $c_{\text{threshold}}=0.9$ for all evaluations.
Iteration Depth. We train a 1.7B TaH-3 with a maximum of three iterations. TaH-3 yields 5.8% average gain over Standard, and 0.8% over TaH-2, as detailed in Appendix A.2.5.
6.4 Behavior Analysis
Token Alternation Patterns. We analyze which tokens TaH selects for deeper iteration. On the validation set, But and So emerge as top candidates, with iteration probabilities of 34% and 18%, respectively. These tokens signal critical contrasting or causal relationships, confirming that models benefit from additional processing at logically complex junctures. Figure 3(a) illustrates how TaH alternates predictions after iteration at these key tokens, suggesting logic refinement behavior. See Appendix A.3.4 for detailed analysis.
Attention Pattern. We visualize the attention pattern of TaH in Figure 10 and Appendix A.3.6. The duo-causal attention learns to focus on different iterations in different heads, extracting broader contexts from multiple depths.
(a) Next-token prediction changes across iterations. Top-2 tokens that think-twice most frequently are visualized.
(b) Distribution of token prediction accuracy across iterations for Ouro and TaH.
Figure 3: Behavior analysis visualizations.
Accuracy Landscape. Figure 3 shows NTP accuracy across iterations. Because Ouro trains all iterations to predict all tokens, predictable tokens across depths largely overlap, leaving more tokens unpredictable by any iteration. TaH instead specializes deeper iterations for hard tokens, reducing overlap and improving coverage under selective iteration.
7 Conclusion
We present TaH, a selective latent iteration method that simultaneously improves model performance and efficiency. Architecturally, TaH introduces duo-causal attention, depth-specific LoRA, and a neural iteration decider, enabling effective selective iteration. An oracle policy guides the stable two-stage training of the tightly coupled LLM backbone and decider. Across nine benchmarks, TaH improves accuracy by 5.3–6.2% over strong baselines with minimal overhead, establishing a new direction for improved reasoning under parameter constraints.
Acknowledgments
This work was supported by National Natural Science Foundation of China (No. 62506197, 62325405, 62104128, U19B2019, U21B2031, 61832007, 62204164, 92364201), Tsinghua EE Xilinx AI Research Fund, and Beijing National Research Center for Information Science and Technology (BNRist).
Ethics Statement
This study raises no ethical issues, it did not involve human subjects or sensitive personal data.
Reproducibility Statement
This paper provides sufficient information to reproduce the reported results. All experiments were conducted using publicly available datasets together with open-source models and code. Appendix A details implementation aspects, including data selection, hyperparameters, and training procedures. To facilitate full reproducibility, we release the code, configuration files, and model checkpoints at https://github.com/thu-nics/TaH.
Appendix A Appendix
A.1 Additional Experiment Setups
A.1.1 Training Recipe
We follow the official training setup of Open-R1 (Hugging Face, 2025) and use the Mixture-of-Thoughts dataset. For our main performance experiments (Section 6.2), we filter samples with output length exceeding 8K tokens from each category (math, code, and QA), then randomly sample 33K examples from each to form a balanced 100K training set. The filtered dataset contains 480M tokens, with 1% reserved for validation. For design choice exploration (Section 6.3), we use only the math subset filtered by 8K output length, resulting in 75K training samples. For 0.6B and 1.7B models, we use a maximum sequence length of 8192 tokens for all methods. For 4B models, we reduce the maximum sequence length to 4096 for Standard, TaH, TaH+, and 3072 for Ouro due to memory constraints. Detailed training hyperparameters are listed in Table 3.
Table 3: Training hyperparameters.
Hyperparameter Value
learning rate 4e-5
max grad norm 0.2
training epochs 5
global batch size 128
warmup ratio 0.03
lr scheduler cosine (min-lr ratio 0.1)
precision bfloat16
A.1.2 Baseline Setups
AlwaysThink. AlwaysThink uses the same architecture as TaH, except that it iterates twice at every token position and uses standard causal attention (attending only to the current iteration depth) instead of duo-causal attention.
Ouro. We implement Ouro architecture following its official design (Zhu et al., 2025), but fine-tune it on the Open-R1 dataset from Qwen-Base initialization, to align with all other methods. During fine-tuning, we adopt the entropy-regularized loss of Ouro. Since the original Ouro is trained from scratch on different data distributions, performance may differ from the original paper.
A.1.3 Parameter Breakdown
Table 4 reports the parameter breakdown of the Standard, TaH, and TaH+ methods. To offset the additional parameters introduced by TaH through LoRA and the iteration decider, we remove one layer from the LLM backbone, ensuring a fair comparison. In practical deployments, we recommend TaH+, which adds only about 3% additional parameters.
Table 4: Parameter breakdown of Standard, TaH, and TaH+. Counts are reported using M (million) and B (billion).
Param. Method Backbone LoRA Iter. Decider Total
0.6B Standard 596M 596M
TaH 580M 10M 5M 595M
TaH+ 596M 10M 5M 611M
1.7B Standard 1.72B 1.72B
TaH 1.67B 17M 18M 1.71B
TaH+ 1.72B 17M 18M 1.76B
4B Standard 4.02B 4.02B
TaH 3.92B 32M 70M 4.02B
TaH+ 4.02B 33M 70M 4.12B
A.1.4 Latent Overthinking Analysis Setup
We set up the oracle experiment in Table 14 to estimate the performance upper bound of our method. The oracle employs the DeepSeek-R1-Distill-Qwen-32B model as a dynamic label generator, replacing the MLP-based iteration decider. During each iteration, we compare the token predictions from the label generator with those from the TaH model. The model continues to iterate only when the top-1 predictions of these two models differ. Due to resource constraints and computational overhead, we evaluated the accuracy only on the first 100 samples from the MATH500 dataset, denoted as MATH100 throughout the paper.
A.2 Additional Experimental Results
A.2.1 Performance on Math Benchmarks
Table 5: Accuracy comparison of different baselines across five benchmarks and two model sizes. Subscripts indicate improvement over Standard. The top two scores for each task and model size are highlighted in bold.
Method
Param. Benchmark Standard Routing SoftThink AlwaysThink TaH TaH+
0.6B AIME25 4.2 1.0 2.5 1.5 4.2 5.0
OlympiadBench 18.8 7.4 19.4 10.2 23.9 24.0
AMC23 23.4 10.9 24.1 15.6 32.5 30.6
MATH500 47.2 27.3 48.8 32.8 51.2 54.2
GSM8K 62.5 45.6 61.3 54.6 64.4 68.8
Average 31.2 18.5 31.2 22.9 35.2/+/+4.0 36.5/+/+5.3
1.7B AIME25 13.3 10.2 12.9 10.0 17.9 14.6
OlympiadBench 33.0 30.6 33.4 30.0 38.8 41.2
AMC23 42.2 42.2 43.1 42.5 48.4 51.2
MATH500 68.4 60.0 68.8 61.8 74.4 73.0
GSM8K 82.1 71.2 79.6 79.3 84.5 85.8
Average 47.8 36.8 47.6 44.7 52.8/+/+5.0 53.2/+/+5.4
4B AIME25 23.3 22.5 22.5 OOM 30.4 28.3
OlympiadBench 47.7 45.0 50.1 50.5 52.0
AMC23 62.8 60.9 64.1 70.3 70.6
MATH500 82.8 76.1 83.2 84.4 85.6
GSM8K 90.5 85.3 90.9 90.4 91.5
Average 61.4 58.0 62.2 65.2/+/+3.8 65.6/+/+4.2
Math Benchmark Evaluation. In this section, all models are trained on the math subset of Open-R1 and evaluated on math benchmarks (see Section A.1.1 for training details). We also add a Routing baseline, which selects a model from a candidate pair for each question. In our experiments, we use two pairs: (1) MobileLLM-R1-360M, Qwen3-1.7B, and (2) Qwen3-0.6B, Qwen3-4B. All candidate models are SFT-trained under the same settings. For each pair, the routing ratio is calibrated so that the average active parameter count matches our 0.6B and 1.7B targets, respectively. Table 5 reports results for 0.6B, 1.7B, and 4B backbones across five challenging math benchmarks. Even with strong Qwen3-Base initialization, existing approaches show limited effectiveness: AlwaysThink and routing methods fail to consistently outperform the standard baseline, while SoftThink yields only marginal gains. In contrast, TaH achieves stable improvements, with average gains of 4.0% (0.6B), 5.0% (1.7B), and 3.8% (4B) over Standard. TaH+, which adds less than 3% parameters, further improves to 5.3%, 5.4%, and 4.2%, respectively. For 0.6B and 1.7B, TaH achieves 8.1–11.3% gains over AlwaysThink, and TaH+ achieves 8.5–12.6% gains. AlwaysThink-4B is not evaluated due to out-of-memory during training.
Out-Of-Distribution (OOD) Performance. We further evaluated the zero-shot generalization capability of models trained solely on math datasets from the main paper. As shown in Table 6, TaH+ demonstrates consistent improvements not only on in-domain math benchmarks (MATH500, AMC23) but also on out-of-domain tasks like MMLU-STEM. This indicates that the thinking patterns learned by TaH+ on math problems are robust and transferrable to broader scientific reasoning tasks.
Table 6: Performance of math-only trained models (0.6B and 1.7B) on in-domain math benchmarks and the out-of-domain STEM benchmark (MMLU-STEM).
Param. Benchmark Standard SoftThink AlwaysThink TaH+
0.6 MATH500 47.2 48.8 32.8 54.2
AMC23 23.4 24.1 15.6 30.6
MMLU-STEM 51.6 51.4 42.6 56.3
Average 40.7 41.4 30.3 47.0
1.7 MATH500 68.4 68.8 61.8 73.0
AMC23 42.2 43.1 42.5 51.2
MMLU-STEM 70.8 70.6 63.8 73.7
Average 60.5 60.8 56.0 66.0
A.2.2 Real-world Efficiency
Setup. We investigate the real-world efficiency of different 1.7B models under our current implementation. All measurements were obtained on a single A800 GPU with a batch size of 1 and a maximum output length of 8192 tokens, using a challenging AIME25 problem where all three methods reached the token limit. Memory usage was profiled using torch.cuda.memory._record_memory_history.
Latency Breakdown. We report the decoding latency, throughput, and a detailed time breakdown for Standard, AlwaysThink, and TaH in Table 7. Here, Iter-1 forward and Iter-2 forward denote the total forward-pass time spent on the first and second latent iterations, respectively; Iter decider is the time for the iteration decider network to judge whether to continue iterating or verbalize; LoRA switching is the overhead of switching LoRA adapters; and Other includes tensor initialization, concatenation, and related bookkeeping.
Table 7: Per-component latency breakdown on a single A800 GPU.
Component Standard TaH AlwaysThink
Latency (s) Ratio(%) Latency (s) Ratio(%) Latency (s) Ratio(%)
Iter-1 Forward 210.6 100.0 229.8 76.2 224.1 30.0
Iter-2 Forward 29.6 9.8 384.7 51.5
Iter. Decider 10.5 3.5
LoRA Switching 7.5 2.5 91.1 12.2
Other 24.1 8.0 47.4 6.3
[…]
Table 8: Real-world decoding performance on a single A800 GPU, tested on AIME25 with 8K max token length. TPS denotes tokens per second.
Standard AlwaysThink TaH
Avg. Depth 1.00 2.00 1.06
Memory (GB) 4.3 6.8 4.6
Latency (s) 210.6 747.2 301.4
TPS 38.9 11.0 27.2
Discussion. We note that our current implementation is not yet optimized at the system level, so there remains room for further efficiency improvements. For example, the LoRA Switching and Other overheads (bookkeeping) are relatively high due to the Python-level implementation of dynamic control flow. These engineering optimizations are largely orthogonal to the algorithmic design of TaH, and we plan to continue refining the implementation to further reduce latency and memory overhead. The theoretical FLOPs and memory access analysis of TaH are provided in Appendix A.2.3.
A.2.3 Theoretical Efficiency Analysis
Following prior work Hoffmann et al. (2022); Yang et al. (2024); Ma et al. (2025), we analyze the computational and memory access overhead of TaH relative to the Standard and AlwaysThink baselines. Table 9 presents the average number of input/output tokens and latent iterations per token across five benchmarks. We use these statistics to calculate the theoretical computation and memory access costs for each method.
[…]
Table 9: Input tokens (shared across methods) and output token / iteration statistics for Standard, AlwaysThink, TaH, and TaH+ (Math-trained version).
Standard AlwaysThink TaH TaH+
Param. Dataset In. Out. Iter. Out. Iter. Out. Iter. Out. Iter.
0.6B AIME25 159 7450 1.00 7316 2.00 7648 1.05 7486 1.06
OlympiadBench 100 6599 1.00 6622 2.00 6631 1.09 6513 1.06
AMC23 85 6377 1.00 6368 2.00 6242 1.05 6145 1.05
MATH500 71 4823 1.00 5350 2.00 4877 1.05 4793 1.06
GSM8K 61 1955 1.00 2844 2.00 1923 1.07 1791 1.07
[…]
1.7B AIME25 159 7195 1.00 7173 2.00 7496 1.06 7498 1.06
OlympiadBench 100 6008 1.00 6484 2.00 6387 1.06 6258 1.06
AMC23 85 5681 1.00 7543 2.00 6122 1.04 5852 1.06
MATH500 71 4004 1.00 4414 2.00 4233 1.06 4286 1.06
GSM8K 61 1451 1.00 1644 2.00 1721 1.08 1686 1.08
[…]
Table 10: Decoding computation (GFLOPs) and memory access (GB) per output token for Standard, AlwaysThink, TaH and TaH+ methods.
Standard AlwaysThink TaH TaH+
Param. Dataset Comp. Mem. Comp. Mem. Comp. Mem. Comp. Mem.
0.6B AIME25 1.47 1.38 3.35 3.14 1.52 1.43 1.57 1.47
OlympiadBench 1.41 1.32 3.21 3.02 1.51 1.42 1.50 1.41
AMC23 1.40 1.31 3.17 2.97 1.43 1.34 1.46 1.37
MATH500 1.31 1.22 2.98 2.80 1.35 1.26 1.39 1.31
GSM8K 1.14 1.06 2.54 2.38 1.19 1.12 1.22 1.14
Average ratio $\times 1.00$$\times$ $\times 2.27$$\times$ $\times 1.04$$\times$ $\times 1.06$$\times$
1.7B AIME25 4.31 4.03 9.45 8.83 4.51 4.21 4.64 4.34
OlympiadBench 4.16 3.88 9.18 8.58 4.36 4.07 4.48 4.18
AMC23 4.12 3.85 9.54 8.91 4.24 3.96 4.43 4.13
MATH500 3.92 3.66 8.45 7.89 4.10 3.83 4.23 3.95
GSM8K 3.62 3.38 7.48 6.98 3.87 3.61 3.98 3.72
Average ratio $\times 1.00$$\times$ $\times 2.19$$\times$ $\times 1.05$$\times$ $\times 1.08$$\times$
A.2.4 Training Dynamics and Continuation Threshold
Figure 4(a) shows the stage 1 training dynamics, and Figure 4(b) reports GSM8K accuracy under different continuation thresholds (with iterate-twice ratios).
(a) Training dynamics of the LLM backbone on Qwen3-0.6B-Base. TaH converges rapidly and achieves lower perplexity.
(b) GSM8K accuracy with respect to continuation threshold. Numbers in brackets show the percentage of iterate-twice tokens.
Figure 4: Training dynamics and continuation threshold analysis.
A.2.5 Iteration Depth Beyond Two
Hard Token Labeling. Previous works have proposed many methods to evaluate the hardness of each tokens in the training data, like through excess loss (Lin et al., 2024; Xie et al., 2023), entropy (Wang et al., 2025b; Chen et al., 2023) and prediction difference (Fu et al., 2025).
For shallow iteration budgets within two ($D_{\max}\leq 2$), we adopt the prediction difference policy. It simply labels the tokens that do not yield top-1 in next-token prediction at the first iteration as hard tokens. Formally, we use a binary halting rule:
$H_{i}^{\pi}\;=\;\begin{cases}0,&\text{if }h_{i}=0\quad\text{(easy token)}\\ D_{\max},&\text{if }h_{i}=1\quad\text{(hard token)}\end{cases}$ (8)
If the iteration depth goes beyond 2 ($D_{\max}>2$), we use the reference model’s cross-entropy as a non-binary indicator of difficulty. Define
$\ell_{i}^{\text{ref}}\;=\;-\log p_{i,\text{ref}}^{(0)}\!\big(t_{i+1}\big).$
We then map difficulty to halting depth via monotone quantile binning:
$H_{i}^{\pi}\;=\;\left\lfloor\mathrm{QuantileRank}\!\big(\ell_{i}^{\text{ref}}\big)\cdot D_{\max}\right\rfloor,$ (9)
where $\mathrm{QuantileRank}(\cdot)\in[0,1]$ is the empirical CDF over the training set (higher loss $\Rightarrow$ deeper halting). This induces per-depth continuation labels $c_{i}^{(d)}=\mathbb{1}[d<H_{i}^{\pi}]$ for $d\in\{0,1,\ldots,D_{\max}\}$.
Experiment Result. Specifically, we train a 1.7B TaH with a maximum per-token iteration count of 3, using oracle labels generated by the method described above. As shown in Table 11, TaH-3 achieves a further improvement of 0.8% on average over TaH-2.
Table 11: Performance comparison between TaH-2 and TaH-3 (maximum per-token iterations of 2 and 3, respectively). Iter.2 and Iter.3 denote the per-token percentages executing two and three iterations, respectively.
Standard TaH-2 TaH-3
Param. Dataset Acc. Acc. Iter.2 Acc. Iter.2 Iter.3
1.7B MATH500 68.4 74.4 5.6 72.6 5.3 0.2
GSM8K 82.1 84.5 7.5 84.8 7.6 0.3
AMC23 42.2 48.4 4.2 48.7 5.1 0.1
OlympiadBench 33.0 38.8 5.7 41.6 5.4 0.2
AIME25 13.3 17.9 6.0 20.4 5.3 0.1
Average 47.8 52.8 5.8 53.6 5.7 0.2
A.2.6 Additional Latent Thinking Methods
Some latent thinking methods require pre-training and use base models other than Qwen3. We also compare with these methods, including Ponder (Zeng et al., 2025). Specifically, we adopt the released pretrained PonderingPythia-1.4B as the base model and perform SFT on the same training data. We observe that the fine-tuned model learns the stylistic patterns of the training data, but still underperforms substantially, which may be attributable to the limited capability of the PonderingPythia-1.4B backbone.
Table 12: Performance on MATH500 and GSM8K-500 (first 500 GSM8K samples)
Method
Dataset Standard-0.6B Ponder-1.4B
MATH500 47.2 2.0
GSM8K-500 62.8 1.8
Avg. 55 1.9
A.2.7 Design Choice Exploration
We conduct controlled ablations where each variant changes a single component from the default TaH configuration to isolate its effect. Table 13 reports the final downstream results, and Figure 12 expands it by showing validation perplexity dynamics across different supervision signals and iteration policies. The naming convention matches Table 13. TaH with token-only supervision and the oracle policy yields lower perplexity than iter. decider and token+latent. Although the dynamic policy achieves the lowest perplexity, it fails on downstream tasks and often produces infinite-loop generations.
Table 13: Ablation study on design choices of TaH-0.6B. Each row varies one aspect from TaH configuration (marked with gray).
Variant MATH500 AMC23 Olympiad Average
TaH 51.2 32.5 23.9 35.9
Model Architecture
Iteration Depth (TaH: Neural decider)
Always-1 47.2 23.4 18.8 29.8//-6.1
Always-2 32.8 15.6 10.2 19.5//-16.4
Attention (TaH: Duo-causal)
Causal@iter1 47.8 24.4 19.4 30.5//-5.4
Causal@current 42.0 23.8 16.4 27.4//-8.5
Depth Adapters (TaH: LoRA + Residual)
w/o LoRA 51.6 29.7 22.4 34.6//-1.3
w/o LoRA & Res. 49.2 22.5 21.2 31.0//-4.9
Training Scheme
Superv. (TaH: Token-only)
Token+latent 49.4 29.6 15.9 31.6//-4.3
Iter. Policy (TaH: Oracle policy)
Decider-based 44.8 24.1 17.3 28.7//-7.2
Dynamic 11.0 2.8 2.7 5.5//-30.4
Discrepancy metric in $\pi$ (TaH: Top1 mismatch)
Cross-entropy 47.4 21.2 20.4 29.7//-6.2
Entropy 42.0 21.9 16.9 26.9//-9.0
Table 14: Impact of iteration schemes on Qwen3-0.6B (first 100 MATH500 samples).
Training Inference Accuracy
Standard Standard 52
AlwaysThink AlwaysThink 38//-14
AlwaysThink TaH-Oracle 77/+/+25
TaH-Oracle TaH-Decider 54/+/+02
TaH-Oracle TaH-Oracle 80/+/+28
A.2.8 Latent Overthinking
To analyze latent thinking patterns, we verbalize tokens from all iteration depths using their last-layer hidden states. The oracle method uses the oracle policy $\pi$ from Section 5.3 for iteration decision. (1) Generation. Since ground-truth tokens are unavailable during generation, we use predictions from the stronger DeepSeek-R1-Distill-Qwen-32B model (Guo et al., 2025) as proxy labels. Table 14 shows that the oracle policy substantially improves performance by verbalizing correct predictions immediately while iterating only on incorrect ones. With our trained iteration decider approximating the oracle, TaH outperforms both Standard and AlwaysThink baselines. However, the ideal oracle policy achieves even higher gains, indicating future potential. (2) Next-token prediction. We evaluate next-token prediction accuracy on the Open-R1 dataset, using the test model itself as the reference model in $\pi$. Figure 1 reveals that AlwaysThink produces more incorrect than correct revisions, demonstrating latent overthinking. In contrast, oracle-controlled iterations substantially increase correct revisions by selectively targeting hard tokens.
A.3 Additional Analysis
A.3.1 Oracle Policy and Hard Token Analysis
Metrics for Hard Token Labeling. We investigate different metrics for defining hard tokens to validate our choice of top-1 prediction mismatch. We compare three labeling strategies:
  1. 1.
    Top-1 Mismatch (TaH Default): Labels a token as hard if the reference model’s greedy prediction differs from the ground truth.
  2. 2.
    Entropy (TaH-Entropy): Labels a token as hard if the reference model’s prediction entropy exceeds a threshold.
  3. 3.
    Cross-Entropy (TaH-CE): Labels a token as hard if the reference model’s cross-entropy loss exceeds a threshold.
To ensure a fair comparison, for TaH-Entropy and TaH-CE, we set the thresholds such that the number of hard tokens in each sample matches the total ratio from the default Top-1 Mismatch policy. This isolates the impact of which tokens are selected, rather than how many.
Figure 5 compares the validation loss, and Table 15 reports downstream accuracy on 0.6B models. While cross-entropy (TaH-CE) improves over entropy labeling (TaH-Entropy), the Top-1 Mismatch policy (TaH) achieves superior performance across all benchmarks. This empirically verifies that directly targeting tokens where the model’s top-1 prediction is wrong is the most effective way to identify hard tokens for TaH.
Figure 5: Validation loss curves of 0.6B models trained with different oracle labeling metrics. The default Top-1 Mismatch policy yields the lowest validation loss.
Table 15: Performance comparison of different difficulty metrics (Entropy, Cross-Entropy, and Top-1 Accuracy) on 0.6B models. All methods mark the same total number of tokens as ”hard.”
Method MATH500 AMC23 OlympiadBench Average
TaH-Entropy 42.0 21.9 16.9 26.9
TaH-CE 47.4 21.2 20.4 29.7
TaH 51.2 32.5 23.9 35.9
Labeling Robustness. We investigate the robustness of hard-token labels with respect to the choice of reference model. We do so by analyzing the consistency of hard-token identification across different model scales (e.g., Qwen3-0.6B, 1.7B, and 4B).
First, we quantify the agreement between models. As shown in Figure 6, hard tokens exhibit high consistency across scales. Notably, even a smaller, less accurate reference model (1.7B) successfully identifies 81% of the hard tokens for a larger model (4B).
Second, to understand the quality of this overlap, we partition tokens into an overlap set (marked as hard by both models) and a non-overlap set (marked as hard by only one model). We plot the cross-entropy loss under each reference model in Figure 7. We observe that overlap tokens have substantially higher average cross-entropy ($\approx 2.0\times$) than non-overlap tokens for all reference models. This indicates that either reference model can identify this core set of ”hard” tokens, which corresponds to positions of genuine, high uncertainty. It reveals a consensus on hardness among models even of different sizes.
(a) Overlap between Qwen3-0.6B and 1.7B
(b) Overlap between Qwen3-1.7B and 4B
Figure 6: Venn diagrams illustrating the overlap of hard-token labels between different reference models. The high overlap proportions indicate that ”hard” tokens are largely consistent across model scales.
(a) Qwen3-1.7B: cross-entropy of overlap vs. non-overlap hard tokens (w.r.t. Qwen3-0.6B).
(b) Qwen3-0.6B: cross-entropy of overlap vs. non-overlap hard tokens (w.r.t. Qwen3-1.7B).
(c) Qwen3-4B: cross-entropy of overlap vs. non-overlap hard tokens (w.r.t. Qwen3-1.7B).
(d) Qwen3-1.7B: cross-entropy of overlap vs. non-overlap hard tokens (w.r.t. Qwen3-4B).
Figure 7: Token-level cross-entropy distributions of overlap and non-overlap hard tokens across different reference model pairs. For each pair of reference models (e.g., Qwen3-1.7B and Qwen3-0.6B), we plot the cross-entropy of tokens labeled as hard by both models (overlap) and by only one model (non-overlap) on both reference models.
A.3.2 Iteration Decider Robustness
We evaluate the iteration decider, trained on the general OpenR1 corpus, across three validation subsets (Math, Code, and QA) to quantify its robustness and cross-domain generalizability. As summarized in Table 16, the decider maintains high decision accuracy across all domains without any retraining.
Table 16: Iteration decider behavior and downstream gains on different validation subsets. The decider is trained once on general OpenR1 and evaluated without retraining.
Metric Math Code QA
Iteration Percentage 7.8% 10.7% 26.6%
Iteration Accuracy 86.7% 82.3% 76.6%
Benchmark Gain over Standard +6.8% +7.9% +5.8%
Despite being invoked on only 7.8-26.6% of tokens, the decider consistently yields 5.8-7.9% absolute accuracy gains over the standard single-pass baseline on all three domains. Moreover, the decider automatically adjusts its iteration rate according to task difficulty: it iterates more frequently on QA (26.6%) than on Math (7.8%), even under a fixed threshold $c_{\text{threshold}}=0.9$. This behavior indicates that the decider responds to intrinsic uncertainty signals in the model’s predictive distribution rather than memorizing domain-specific patterns, consistent with the token-level difficulty analysis in Appendix A.3.3.
A.3.3 Hard Token Identifiability
(a) Math
(b) Science
(c) Code
Figure 8: Output Logit entropy distribution at the first iteration of TaH, categorized by oracle policy’s difficulty labels (hard token) on the OpenR1 validation set (Math, QA, Code). The distinct separation between distributions confirms that TaH’s internal logits provide a strong, task-agnostic signal for identifying hard tokens.
Why is the iteration decider robust and generalizable across tasks? We investigate this by analyzing the intrinsic properties of ”hard” tokens.
We compute the token entropy of hard and easy tokens across three diverse subsets of the OpenR1 dataset (Math, Science, and Code). As shown in Figure 8, hard tokens exhibit a universal signature of significantly higher entropy ($5\times$) compared to easy tokens. This distinct separation confirms that ”hardness” is an intrinsic, robustly identifiable property of the model’s predictive state, rather than a complex, task-specific pattern. Given this clear signal, the neural iteration decider can easily learn reliable classification strategies that generalize well across different domains.
A.3.4 Token Alternation Pattern
We analyze tokens that most frequently trigger a second iteration (”think-twice” tokens). For each token type $t$, we compute the continuation rate
$\Pr\big(c_{i}^{(1)}>c_{\text{threshold}}\mid t_{i}=t\big),$
using the inference threshold $c_{\text{threshold}}=0.9$ (Section 5.3). We estimate this quantity on the Open-R1 validation set and, for diagnostics, randomly sample 10K token positions ($\approx$0.4% of tokens) to track whether the next-token prediction switches between depth 1 and depth 2. This setting quantifies which token types most often trigger an additional iteration and how often iteration alters the predicted next token.
Table 17: Conditional probabilities of continuation confidence and next-token distribution.
Token $T_{1}$ $P(c^{(1)}>c_{\text{threshold}}\mid t^{(1)}=T_{1})$ Token $T_{2}$ $P(t^{(2)}=T_{2}\mid t^{(1)}=T_{1})$
But 34.3% So 13.63%
Wait 12.17%
Therefore 8.95%
So 17.7% So 28.17%
Therefore 13.67%
But 4.89%
A.3.5 Iteration Decision Error
Figure 9: Iteration-decider accuracy vs. epoch (Qwen3-0.6B).
We analyze how iteration decision accuracy affects TaH’s end-to-end response quality, since iteration decider will not be perfect as shown in Figure 9. To this end, we randomly inject errors into the oracle iteration-decider predictions at different rates. Formally, we denote the original oracle prediction as the label $l\in\{0,1\}$ and the altered prediction as the output $o\in\{0,1\}$. We define the iter. error as the total proportion of deliberately introduced errors:
$\text{iter. error}=P(l\neq o)=\underbrace{P(l=1,o=0)}_{\textit{underthink rate}}+\underbrace{P(l=0,o=1)}_{\textit{overthink rate}}.$ (10)
We further distinguish the impacts of overthinking and underthinking. Here, overthinking refers to cases where the decider incorrectly signals continue, while underthinking corresponds to cases where it incorrectly signals stop. Table 18 shows how TaH’s MATH100 accuracy varies with different iteration error rates. We quantify these effects by fitting a linear model to the data:
$\text{accuracy}=-1.41\times\text{underthink rate}-2.73\times\text{overthink rate}+0.81.$
This analysis indicates that inaccurate iteration decisions are the main factor behind the performance gap between TaH and its oracle variant, with overthinking being the dominant source of performance gaps.
Table 18: TaH performance under different iteration-decider error rates. All values are reported in percentages.
Iter. Error (%) Underthink (%) Overthink (%) MATH100 Accuracy (%)
0.0 0.0 0.0 80.0
2.8 2.8 0.0 78.0
10.0 1.5 8.5 55.4
15.0 2.1 12.9 45.2
20.0 2.5 17.5 27.1
22.1 0.0 22.1 21.6
A.3.6 Duo-causal Attention Pattern
We perform forward computation on 100 samples, each with a length of 128 tokens. Figure 10 shows the average attention weights of three representative attention heads in the second iteration of the TaH model. The left panel illustrates a head that mainly attends to keys from the first iteration. The middle panel shows a head focusing on keys from the second iteration. The right panel displays a head with a balanced attention distribution. These results suggest that the TaH model, under the duo-causal attention mechanism, can automatically learn diverse attention patterns across layers and heads.
Figure 10: TaH duo-causal attention pattern.
Figure 12 further presents the total attention scores assigned to keys in the first iteration. It can be seen that the first layer tends to focus more on keys from the second iteration. Different layers also exhibit varying attention behaviors.
Figure 11: TaH mean and standard deviation of attention weights (key from iteration 1) across layers in iteration 2.
Figure 12: Validation perplexity for different training schemes.
A.4 Implementation Details
A.4.1 Duo-causal attention implementation
[…]
Figure 13 illustrates the implementation of duo-causal attention, with the formal definitions provided below.
(1) KV cache concatenation. At depth $d$, we form the visible K/V sequence by concatenating all shallower-to-current depths along the sequence dimension:
$\mathrm{KV}^{(\leq d)}\,=\,[\,\mathrm{KV}^{(1)}\,;\,\mathrm{KV}^{(2)}\,;\,\cdots\,;\,\mathrm{KV}^{(d)}\,].$
This realizes the accessible set in Equation 6, allowing deeper iterations to access all shallower iterations while preserving positional causality. The KV cache is managed by iteration depth during decoding, as shown in Figure 13(b). The fragmented KV-cache management strategy is standard in existing LLM serving systems (Kwon et al., 2023; Zheng et al., 2024).
(2) Two-dimensional causal mask. For a query $(i,d)$, a key $(j,k)$ is attendable iff $j\leq i$ and $k\leq d$. We implement this as an additive attention mask with 0 for allowed entries and -$\infty$ otherwise, enforcing positional and iteration causality jointly. Figure 13(c) visualize the landscape of the duo-causal attention mask. When $d=1$ for all tokens, the rule reduces to standard causal attention.
(3) Compatibility with efficient attention. The mask is provided in the standard additive form and the concatenated K/V remain contiguous along the sequence dimension, matching the usual scaled dot-product attention interface. As a result, duo-causal attention is directly compatible with optimized kernels such as FlashAttention, without kernel modifications.
A.5 Additional Related Work
Signal-guided Control. These methods keep reasoning in token space but steer computation by inserting control tokens. They add filler tokens (e.g., dots) (Pfau et al., 2024), learnable [PAUSE] tokens (Goyal et al., 2024; Kim et al., 2025) for extra compute during decoding. They are lightweight but constrained to discrete-token interventions with limited latent control.
Latent Optimization. These methods perform autoregressive reasoning directly in internal representations, emitting little or no intermediate text. They distill CoT into continuous embeddings via progressive replacement (Hao et al., 2024; Cheng & Van Durme, 2024), hidden-state alignment (Su et al., 2025; Liu et al., 2024), or logit-weighted embeddings (Zhang et al., 2025). While efficient, these methods sacrifice interpretability, with training-based ones requiring heavy mitigation from verbal LLMs.
Instead of using the shared model parameter multiple times through latent iteration, previous work also proposes layer skipping methods for dynamic computing allocation.
Layer Skipping. Layer skipping aims to accelerate LLM inference by dynamically bypassing certain layers for specific tokens. Some methods use a learnable module to make real-time skipping decisions. MoD (Raposo et al., 2024) uses a top-k router to select a subset of tokens for processing, while FlexiDepth (Luo et al., 2025) uses a plug-in router to determine whether a layer should be bypassed. Others use a fixed strategy to skip layers. SkipDecode (Del Corro et al., 2023) enforces a monotonically decreasing number of active layers during generation. However, these methods still require loading the entire model’s parameters, resulting in a large memory access overhead. Instead of skipping some layers, TaH adds computational depth by allowing core tokens to undergo multiple refinement iterations. This approach provides greater computational depth without increasing the model’s parameter count.
A.6 Limitations and Future Work
Comparison with Official Qwen3 Models. Official Qwen3 models are trained on different data distributions and scales, and use different training procedures, including on-policy distillation (Yang et al., 2025). By contrast, our models use SFT only on limited, publicly accessible data. Consequently, performance may differ between the two.
Future Work. We consider advanced training techniques largely orthogonal to TaH. Promising directions include: (1) integrating TaH with online distillation to provide stronger supervision; (2) applying reinforcement learning to optimize the compute–accuracy trade-off and to learn continuation policies beyond imitation. Due to resource constraints, this work focuses on SFT with public data; we leave these combinations for future work.
A.7 Use of Large Language Models (LLMs)
Large language models (e.g., GPT-5, Gemini) were used exclusively for proofreading and minor copyediting. They were not employed to generate scientific content, contribute to research ideation, or design methods. All research ideas, methods, and conclusions are solely the responsibility of the authors.
A.8 Examples
We use an example question from AIME25 and responses from Standard, TaH and AlwaysThink to provide an intuitive understanding of our method. Only TaH can successfully answer this question. Blue-highlighted tokens indicate positions that undergo two iterations.
Text 1. Question On $\triangle ABC$ points $A,D,E$, and $B$ lie that order on side $\overline{AB}$ with $AD=4,DE=16$, and $EB=8$. Points $A,F,G$, and $C$ lie in that order on side $\overline{AC}$ with $AF=13,FG=52$, and $GC=26$. Let $M$ be the reflection of $D$ through $F$, and let NN be the reflection of $G$ through $E$. Quadrilateral DEGFDEGF has area 288. Find the area of heptagon AFNBCEMAFNBCEM.
Text 2. Standard Response Okay, let me try to figure out this problem step by step. It’s about a heptagon formed by reflecting points and finding its area. …We have triangle ABCABC. On side ABAB, points $A,D,E$, and $B$ are in that order. The lengths are given as $AD=4$, $DE=16$, and $EB=8$. So $AB=28$ …On side ACAC, points $A,F,G$, and $C$ are in that order with $AF=13$, $FG=52$, and $GC=26$. So $AC=91$ …Let me place point AA at $(0,0)$ and $B$ at $(28,0)$ …Let $C$ be at $(91,h)$ …Then $D=(4,0)$, $E=(20,0)$, $F=\big(13,\frac{13h}{91}\big)$, $G=\big(65,\frac{65h}{91}\big)$ Area of quadrilateral DEGFDEGF can be calculated …Applying the shoelace formula …Area $\frac{1}{2}\cdot\frac{1248h}{91}=\frac{624h}{91}\;\mathbf{=}\;6h$ So with $288$: $6h=288\Rightarrow h=\frac{288}{6}=48$ Reflections: $M$ is the reflection of $D$ through $F$ and $N$ is the reflection of $G$ through EEM=(22,0)M=(22,0), N=(39,0)N=(-39,0), $B=(28,0)$, $C=(91,48)$, $E=(20,0)$ Using the shoelace formula for AFNBCEMAFNBCEM in order …Sum =0+0+0+1344+(960)+0+0=384=0+0+0+1344+(-960)+0+0=384 Take absolute value and multiply by $1/2$: Area $\frac{1}{2}\cdot|384|=192$ Final Answer $\boxed{192}$
Text 3. TaH Response Okay, let’s try to tackle this geometry problem. It seems a bit complex with all these points and reflections, …First, let me visualize triangle ABC. Points D, E, and B are on side AB with AD = 4, DE = 16, and EB = 8. …Let me set up a coordinate system. Let’s place point A at the origin $(0,0)$ …Then, since ABAB is a line segment, let me assume ABAB is along the $x$-axis. So point $B$ would be at $(28,0)$ …Now, points $F$, G, $C$ are on ACAC …Let me define coordinates for $C$ as $(c_{x},c_{y})$ Compute the sum of $x_{i}y_{i+1}-x_{i+1}y_{i}$ for each $i$ Sum of all terms: $0+\frac{1300c_{y}}{91}+0->\frac{52c_{y}}{91}=\frac{(1300-52)c_{y}}{91}=\frac{1248c_{y}}{91}$ Then, the area is $\frac{1}{2}\cdot\left|\frac{1248c_{y}}{91}\right|=\frac{624c_{y}}{91}$ Compute the area of this heptagon … Using shoelace formula … Term2: 240-\!240  Term3: $840$  Term4: $1176$  Term5: 840-\!840  Term6: $240$ Sum: 0240+840+1176840+240+0=11760-240+840+1176-840+240+0=1176 Take absolute value and multiply by $1/2$: Area $\frac{1}{2}\cdot|1176|=588$ Final Answer $\boxed{588}$
Text 4. AlwaysThink Response Okay, so I need to find the area of the heptagon AFNBCEM. Let me first try to visualize the problem. There’s triangle ABC. On side AB, there are points A, D, E, and B. The order is A, D, E, B. The distances given are AD=4, DE=16, and EB=8. So, starting at A, moving 4 units to D, then 16 units to E, then 8 units to B. So, the total length of AB is AD + DE + EB = 4 + 16 + 8 = 28 units. Got that. Points in order: D(4,0), E(20,0), F(13,h), G(65,h), back to D(4,0). Compute the terms: Sum these terms: 0+20h52h4h=36h0+20h-52h-4h=-36h Take absolute value and multiply by 1/2: Area=(1/2)|36h|=(1/2)36|h|=18|h|Area=(1/2)*|-36h|=(1/2)*36|h|=18|h| But the area is given as 288, so $18|h|=288\Rightarrow|h|=16$. Now, apply the shoelace formula …Therefore, total sum is $8640/17$ Take absolute value and multiply by 1/2: $Area=(1/2)*|8640/17|=8640/34=4320/17\approx 254.1176$ Therefore, the answer is $4320/17$. Final Answer $\boxed{\frac{4320}{117}}$